OpenStudy (vcabral1134):

Help me plz will fan and give medal! Write the explicit formula for the geometric sequence. Then find the fifth term in the sequence. a^1 = −3, a^2 = 15, a^3 = −75 a. an = −3 ⋅ (−5)n ; 9375 c. an = −3 ⋅ (5)n ; –1875 b. an = −5 ⋅ (−3)n − 1 ; –405 d. an = −3 ⋅ (−5)n − 1 ; –1875

Nnesha (nnesha):

\[\huge\rm a_n = a_1 \times r^{n-1}\] equation to find nth term of geometric seq. where a_1 is first term r= ratio n= term that u need

OpenStudy (vcabral1134):

thank you

Nnesha (nnesha):

to find common ratio divide next term with the previous one \[\large\rm r=\frac{ a_2 }{ a_1}~~or~~\frac{ a_3 }{ a_2} \]

Nnesha (nnesha):

a_2 = 2nd term a_3 = 3rd term

OpenStudy (vcabral1134):

r=-5

OpenStudy (vcabral1134):

D @Nnesha

OpenStudy (vcabral1134):

Am I right??

Nnesha (nnesha):

let me check wait a sec

OpenStudy (vcabral1134):

ok

Nnesha (nnesha):

right

OpenStudy (vcabral1134):

Yay!!

Nnesha (nnesha):

yahoo!